Will rate, please help!!!- Use the First Derivative Test to find a) the intervals over which f(x) is increasing and decreasing, and b) the local extre
Answer #1

fors = 122 - 2 1 fx> 4x3x? - 3 x 9 +0 = **-3n = X(4-3) for function to be increasing fresto > n(n-3))0 : xoo and (4-3) > >0,